# Distributionally robust expected shortfall for convex risks

### v1

@@ Abstract
We study distributionally robust expected values under optimal transport distance with a quadratic cost function. In general the duality method, for this computation for the payoff function $f$, requires the computation of the $\lambda c$-transform $f^{\lambda c}$. We show that under the quadratic cost function there exists an intuitive and easily implementable representation of $f^{\lambda c}$, if $f$ is convex and piecewise linear. We apply this to the robust expected shortfall under the risk-neutral measure of an unhedged call option, from the point of view of the writer, as well as that of a portfolio mixing underlying shares with a call and a put option.
@@ 1 Introduction
The study of distributional risk attempts to quantify the “consequences of using the wrong models [or statistical distributions] in risk measurement, pricing and portfolio selection” [4]. Causes of distributional risk include bad choice of stochastic process, wrong dependence assumptions, and parameter drift. Examples of wrong distributional assumptions are well known. The demise of LCTM in 1999 is partly blamed on overreliance on Gaussian logreturns, and the breakdown of credit models in the 2008 financial crisis, is partly blamed on the unexpected behaviour of correlations during crisis periods. Of course, not all model risk [9] is distributional risk. Also, the study of distributional risk is intended here not as a substitute for the correct dynamics, but as an attempt to probe, and possibly compare, the distributional risks of different portfolios, if no further knowledge about the distribution is assumed than that it may be off from the correct distribution by a certain distributional “distance.”
@@
One could assess distributional risk using parametric or non-parametric approaches. Parametric approaches include describing the sensitivity of an option price with regard to volatility and other statistically important parameters. A related approach is uses parameter intervals.
@@
Non-parametric approaches are usually based on the question of how much a financially relevant calculation could be affected if the underlying “baseline” or nominal distribution is varied within an “uncertainty set” of alternative distributions, the set of alternatives not characterized by a few parameters. Typically the uncertainty set comprises those distributions that differ from the baseline by at most a certain amount $\theta>0$. There are several prominent ways to quantify this difference between the baseline and alternative distributions. In quantitative finance. Glasserman and Xu [7] pioneered the use of “relative entropy” as distance measure between distributions. See, for example Feng et al [6] for an application to option pricing models. Another quantification, which is the focus of this paper, is by optimal transport-like distances of measures. Other statistical “divergence” measures, for example those suited to the tails of distributions, are also possible.
@@
In a topic as broad as the relationship between map and reality, the above-mentioned categories are by no means exhaustive. It may also be mentioned that the term “distributionally robust”, which is widespread in the literature, is not meant to suggest robustness with respect to all possible changes in distribution, which is in general an impossible goal. Instead, in the non-parametric approaches referred to above, the distributional robustness is always relative to an ambiguity tolerance parameter $\theta$. For larger values of $\theta$ the calculation is more conservative. The non-robust case corresponds to $\theta=0$.
@@
The optimal transport distance $d_{c}$ to quantify the distance between probability measures on $\mathbb{R}^{d}$, is based on a cost function $c(x,y)$ for $x,y\in\mathbb{R}^{d}$. Section 2 gives an example of such a distance computation for the quadratic cost function. The robust problem is then to maximize $\int fd\nu$ subject to the constraint that $\nu\in\mathcal{N}$, where $\mathcal{N}$ denotes the uncertainty set of probability measures for which $d_{c}(\nu,\nu_{0})\leq\theta$. In this optimization problem the uncertainty set, involving a set of probability distributions or measures, is infinite dimensional. For applications the dual formulation is most useful, see for example [1, Theorem 2.4 and Equation (7)]
$\sup_{\nu\in\mathcal{N}:\ d_{c}(\nu,\nu_{0})\leq\theta}\int fd\nu=\inf_{\lambda\geq 0}\left\{\lambda\theta+\int f^{\lambda c}d\nu_{0}\right\}.$ (1)
@@
Thus the infinite-dimensional problem is reduced to one-dimensional optimization problem, provided that $f^{\lambda c}$ is known. We show in section 2 that in an important particular case, where $f$ is convex and piecewise-linear, there is an easy way of determining $f^{\lambda c}$. In section 3 we derive a representation of robust Expected Shortfall in terms of an optimization over two parameters of an expected value calculation, and apply this to derive an analytical expression for the robust, subject to ambiguity tolerance $\theta$, Expected Shortfall of a call option, from the point of view of the writer, under the risk-neutral distribution. In section 4 we apply this representation to portfolio robust expected shortfall minimization.
@@ 2 $\lambda c$-transform for convex, piecewise-linear, functions
where $\textrm{Cpl}(\mu,\nu)$ denotes the set of “couplings” of $\mu$ and $\nu$; that is, the set of probability measures $\pi$ on $X\times X$ with first marginal distribution equal to $\mu$, and second to $\nu$. A coupling will always exist, take for example $\pi$ as the product measure $\mu\otimes\nu$. In this paper we are only concerned with the quadratic cost function $c(x,y)=\frac{1}{2}\|x-y\|^{2}$ for $x,y\in\mathbb{R}^{d}$, where $\|\cdot\|$ denotes the Euclidean norm of a vector in $\mathbb{R}^{d}$. (The factor $\frac{1}{2}$ is convenient in calculations where the cost function is differentiated.) With the quadratic cost function, it holds [11, Theorem 4.1] that $d_{c}(\mu,\nu)<\infty$ if $\mu$ and $\nu$ have finite second moments; that is, if $\int_{X}\|x_{0}-x\|^{2}d\mu(x)<\infty$ and $\int_{X}\|x_{0}-x\|^{2}d\nu(x)<\infty$ for any $x_{0}\in X$.
@@
As mentioned in the introduction, the robust expected value problem is tractable, if the $\lambda c$-transform $f^{\lambda c}$ is available. In an important special case, there is a simple representation.
@@
Suppose that there exists vectors $m_{i}\in\mathbb{R}^{d}$ and scalars $c_{i},\ 1\leq i\leq n$, so that
$f(x)=\max_{1\leq i\leq n}\left\{\langle m_{i},x\rangle+c_{i}\right\},\ x\in\mathbb{R}^{d}.$
Then
$f^{\lambda c}(x)=\max_{1\leq i\leq n}\left\{\langle m_{i},x\rangle+c_{i}+\frac{1}{2\lambda}\|m_{i}\|^{2}\right\},\ x\in\mathbb{R}^{d}.$
Theorem 1 is proved in the Appendix. A heuristic interpretation, of some features of the formula, is given in Figure 2, where the $\lambda c$-transforms, of payoff functions that differ by a constant $k$, also differ by $k$. This is in line with the intuition that a payoff, independent of the underlying, is not exposed to the distribution risk of the underlying.
@@
Call and put option with same strike. Let $f(x)=\max\{x-K,K-x,0\}.$ Using Theorem 1, $f^{\lambda c}(x)=\max\left\{x-(K-\frac{1}{2\lambda}),(K+\frac{1}{2\lambda})-x,0\right\}$.
@@
Figure 3: $\lambda c$-transform, call and put payoff with same strike.
@@
$f^{\lambda c}(x)=\max_{i=1,\dots,4}\left\{\langle m_{i},c_{i}\rangle+\frac{\|m_{i}\|^{2}}{2\lambda}\right\}.$ (4)
@@ 3 Robust Expected Shortfall
We consider expected shortfall as a risk measure $\rho$ on payoff functions on $\mathbb{R}^{d}$. For risk measures on functions, see [5]. In the distributional robustification process, we associate a risk measure $\rho_{\nu}$ to each $\nu$ in the uncertainty set. In particular, Expected Shortfall of the same payoff under different probability distributions, are seen as different risk measures in the sense of giving different risk assessments. It is easy to see that in this setup the robustification $\tilde{\rho}(f):=\displaystyle\sup_{d_{c}(\nu,\nu_{0})\leq\theta}\rho_{\nu}(f)$ is a coherent risk measure, if each $\rho_{\nu}$ is a coherent risk measure. In particular $\tilde{\rho}(0)=0$.
For simplicity we consider a liability $f$ on an atomless space, so that $\textrm{ES}_{\beta}^{\ \nu}(f)=\frac{1}{1-\beta}\int_{w}^{\infty}f(x)d\nu(x)$, where $w:=\textrm{VaR}_{\beta}^{\nu}(f)$ is the $\beta$-quantile of $f$ under the measure $\nu$, and $\beta$ the confidence level under which Expected Shortfall is computed. Let us denote the baseline risk-neutral measure by $\nu_{0}$. As in [1] we use the “minimizing property” that characterizes $\textrm{ES}_{\beta}$, see Rockafellar and Uryasev [10, Theorem 1],
@@
This expression will be referred to as the dual formula for Expected Shortfall, as it can also be derived using the dual representation of that risk measure. We recall from [10] that $\alpha^{*}:=\textrm{VaR}_{\beta}(f)$ is a minimizer in this expression. In this section $d=1$.
Our uncertainty set is
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be Lipschitz continuous. The robust $\textrm{ES}_{\beta}(f)$ namely
@@
$\tilde{\rho}(f)=\inf\left\{\lambda\theta+\alpha+\frac{1}{1-\beta}\int_{\mathbb{R}}((f(x)-\alpha)^{+})^{\lambda c}\ d\nu_{0}(x):\ \alpha\in\mathbb{R},\lambda\geq 0\right\}.$ (6)
@@
$=$ $\sup_{\nu\in\mathcal{N}}\min_{\alpha\in\mathbb{R}}\left\{\alpha+\frac{1}{1-\beta}\int_{-\infty}^{\infty}(f(x)-\alpha)^{+}\,d\nu(x)\right\}$
$=$ $\sup_{\nu\in\mathcal{N}}\min_{v_{min}\leq\alpha\leq v_{max}}\left\{\alpha+\frac{1}{1-\beta}\int_{-\infty}^{\infty}(f(x)-\alpha)^{+}\,d\nu(x)\right\},$
@@
To see that an upper bound $v_{max}$ exists, let $\|f\|_{Lip}$ be the Lipschitz constant of $f$ and $W_{1},W_{2}=\sqrt{d_{c}}$ denote the respective p-Wasserstein distances. It follows from the Markov inequality, 1-Wasserstein duality [11, Remark 6.5], and the basic property $W_{1}\leq W_{2}$ [11, Remark 6.6] that for $t>0$ ,
$\nu(\{x:f(x)>t\})$ $\leq$ $\frac{1}{t}\int_{\mathbb{R}}|f(x)|d\nu(x)$
$\leq$ $\frac{1}{t}\left(\int_{\mathbb{R}}|f(x)|d\nu_{0}(x)+\|f\|_{Lip}W_{1}(\nu,\nu_{0})\right)$
$\leq$ $\frac{1}{t}\left(\int_{\mathbb{R}}|f(x)|d\nu_{0}(x)+\|f\|_{Lip}\sqrt{\theta}\right)$
@@
Since $\alpha$ ranges over a compact set, and the function $(\alpha,\nu)\mapsto\alpha+\frac{1}{1-\beta}\int_{-\infty}^{\infty}(f(x)-\alpha)^{+}\,d\nu(x)$, being convex in $\alpha$ and linear in $\nu$, is easily seen to be “convex-concavelike” [3], we may use Ky Fan’s minimax theorem. Thus
$\tilde{\rho}(f)$ $=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\sup_{\nu\in\mathcal{N}}\left\{\alpha+\frac{1}{1-\beta}\int_{-\infty}^{\infty}(f(x)-\alpha)^{+}\,d\nu(x)\right\}$
$=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\sup_{\nu\in\mathcal{N}}\int_{-\infty}^{\infty}\alpha+\frac{1}{1-\beta}(f(x)-\alpha)^{+}d\nu(x).$
Applying Equation (1) to the function $x\mapsto\alpha+\frac{1}{1-\beta}(f(x)-\alpha)^{+}$, we get the result. ∎
@@
Since $\nu_{0}$ is a pricing measure, $\int g^{\lambda c}(x)\,d\nu_{0}(x)=\textrm{call}(k+\alpha-\frac{1}{2\lambda})$. Thus
$\tilde{\rho}(f)=\min_{v_{min}\leq\alpha\leq v_{max}}\inf_{\lambda\geq 0}\left\{\lambda\theta+\alpha+\frac{1}{1-\beta}\left[\lambda\theta+\textrm{call}(k+\alpha-\frac{1}{2\lambda})\right]\right\}.$ (7)
@@
Therefore, in contrast to some of the earlier robustifications of ES [12][1, Table 1], the “robustification correction” $\tilde{\rho}(f)-\rho_{\nu_{0}}(f)$ for Expected Shortfall is a function of $f$. Indeed,
$\tilde{\rho}(f)$ $=$ $\rho_{\nu_{0}}(f)+\sqrt{\frac{2\theta}{1-\beta}}\text{ if }k\leq q_{\beta},but$ (9)
$\tilde{\rho}(f)$ $=$ $\rho_{\nu_{0}}(f)+(q_{\beta}-k)+\sqrt{\frac{2\theta}{1-\beta}}\text{ if }k>q_{\beta}.$ (10)
This difference is due to the definition of the risk measure in terms of the payoff instead of the distribution, and the fact that in the ES maximization process, imagining an adversary that wants to maximize risk, it is only worthwhile to transport mass that is already on the right hand side of the strike price.
Notably the high distributional sensitivity of an Expected Shortfall calculation if a confidence level $\beta$ very close to one is computed, is hereby quantified.
4 Robust expected shortfall of a three-asset claim
We consider a liability to pay a claim that is given by a portfolio invested in assets A, B and C of Example 3. (That is, the Expected Shortfall is calculated for an investor that is short in the assets and options.) The baseline distribution on $\mathbb{R}^{3}$ is determined by the independent lognormal distributions of the three shares that comprise asset A, provide the underlying for call options of B, or the put option of C. The lognormal parameters are $(\mu_{A},\mu_{B},\mu_{C})=(0.0601,0.0529,0.0713)$ and $(\sigma_{A},\sigma_{B},\sigma_{C})=(0.1836,0.1198,0.2167)$ are chosen, using the approach of [2], to ensure that assets A, B and C have return means $8\%$ and standard deviation $20\%$. Our confidence level is $\beta=95\%$. The strike prices are $K_{2}=K_{3}=1$.
At the beginning of the time period of one year, one monetary unit is received. Expected Shortfall is calculated on the net liability at the end of the year, which is the portfolio payoff less one, reported as a percentage of the initial receipt.
Expected Shortfall is calculated using Equation 5, with dimension $d=3$ and the payoff function in Equation 3. For simplicity we use the Trapezoidal Rule for the numerical integration. Although greater accuracy can be obtained with higher order quadrature, the value in this calculation of the further decimal accuracy due to such refinement, is debatable because the goal is to have a reference point for comparison. That is, exact values are probably less important than the differences between the values for different portfolios. (Higher order quadrature could reduce computation time significantly, though, because fewer function evaluations would be required.) The function of $\alpha$ that is to be minimized, is convex and we use a straightforward Golden Section search for the minimum.
For the robust Expected Shortfall we use Equation 6. This involves the transformed payoff function of Equation 4, and another layer of minimization, also over a convex function (see [13]), is also done with Golden Section search.
We display the results in Table 1 below. The case $\theta=0$, where $\theta$ is the ambiguity tolerance parameter, corresponds to the non-robust Expected Shortfall, and $\theta=1$ to the robustified calculation.
$w_{1}$ $w_{2}$ $w_{3}$ $\theta$ Robust Expected Shortfall with ambiguity tolerance $\theta$
$1/3$ $1/3$ $1/3$ 0 35%
$1/3$ $1/3$ $1/3$ 1 70%
$0.1$ $0.8$ $0.1$ 0 48%
$0.1$ $0.8$ $0.1$ 1 100%
$0.1$ $0.1$ $0.8$ 0 52%
$0.1$ $0.1$ $0.8$ 1 105%
Table 1: Expected Shortfall for different sets of portfolio weights. Non-robust case is $\theta=0$.
Comparing the difference between the fourth and third rows, or between the sixth and fifth rows, to the difference between the second and first rows, we can quantify the extent to which the exposure to the assets that have option components, are significantly more subject to distributional risk, than the equally weighted exposure.

### v2

# Distributionally robust Expected Shortfall for convex payoffs

@@ Abstract
We study distributionally robust expected shortfall when the distribution of the underlying is perturbed by a size quantified with optimal transport distance based on the quadratic cost function. In the dual version of the robust expectation problem, which is part of the robest expected shortfall problem, the computation of the so-called $\lambda c$-transform $f^{\lambda c}$ of payoff $f$ is required. We show that under the quadratic cost function there exists a tractable representation of $f^{\lambda c}$, if $f$ is convex. Furthermore, we show that robust expected shortfall can be characterized as the solution of a 2-dimensional minimization problem. We apply these results to obtain a closed-form formula for robust, with respect to the risk-neutral distribution, expected shortfall of an unhedged call option, from the point of view of the writer.
@@ 1 Introduction
The study of distributional risk attempts to quantify the “consequences of using the wrong models [or statistical distributions] in risk measurement, pricing and portfolio selection” [BreuerCsiszár2016]. Causes of distributional risk include bad choice of stochastic process, wrong dependence assumptions, and parameter drift. Examples of wrong distributional assumptions are well known. The demise of LCTM in 1999 is partly blamed on overreliance on Gaussian logreturns, and the breakdown of credit models in the 2008 financial crisis, on the unexpected behaviour of correlations during crisis periods. Model risk [7] is broader than distributional. It may also be mentioned that the study of distributional risk is intended here not as a substitute for the correct dynamics, but as an attempt to probe, and possibly compare, the distributional risks of different portfolios, if no further knowledge about the distribution is assumed than that it may be off from the correct distribution by a certain distributional “distance.”
@@
One could assess distributional risk using parametric or non-parametric approaches. Parametric approaches include describing the sensitivity of an option price with regard to volatility and other statistically important parameters. A related approach is to parameter intervals.
@@
Non-parametric approaches are usually based on the question of how much a financially relevant calculation could be affected if the underlying “baseline” or nominal distribution is varied within an “uncertainty set” of alternative distributions, the set of alternatives not characterized by a few parameters. Typically the uncertainty set comprises those distributions that differ from the baseline by at most a certain amount $\theta>0$. There are several prominent ways to quantify this difference between the baseline and alternative distributions. In quantitative finance, Glasserman and Xu [6] pioneered the use of relative entropy as distance measure between distributions. See, for example Feng et al [5] for an application to option pricing models. Another quantification, which is the focus of this paper, is by optimal transport-like distances of measures. Other statistical “divergence” measures, for example those suited to the tails of distributions, are also possible.
@@
The above-mentioned list of categories is of course by no means complete. It may also be mentioned that the term “distributionally robust”, which is widespread in the literature, is not meant to suggest robustness with respect to all possible changes in distribution, which is in general an impossible goal. Instead, in the non-parametric approaches referred to above, the distributional robustness is always relative to an ambiguity tolerance parameter $\theta$. For larger values of $\theta$ the calculation is more conservative. The non-robust case corresponds to $\theta=0$.
The optimal transport distance $d_{c}$ to quantify the distance on the set $P(\mathbb{R}^{d})$ of probability measures on $\mathbb{R}^{d}$, is based on a cost function $c(x,y)$ for $x,y\in\mathbb{R}^{d}$. In this paper $c(x,y)=\frac{1}{2}\|x-y\|^{2}$, called the quadratic cost function, with $\|\|$ denoting the Euclidean norm. The robust expectation problem is
$P:=\sup\left\{\int fd\nu:\ \nu\in\mathcal{N}\right\},$ (1)
@@
where $\mathcal{N}=\{\nu\in P(\mathbb{R}^{d}):\ d_{c}(\nu,\nu_{0})\leq\theta\}$. In this optimization problem the uncertainty set, involving a set of probability distributions or measures, is infinite dimensional. For applications a duality result such as that by Bartl et al [1, Theorem 2.4 and Equation (7)]
$P=D:=\inf_{\lambda\geq 0}\left\{\lambda\theta+\int f^{\lambda c}d\nu_{0}\right\},$ (2)
where $f^{\lambda c}(x)=\sup_{y\in\mathbb{R}^{d}}\{f(y)-\lambda c(x,y)\}$, is usually needed. Thus the infinite-dimensional problem is reduced to one-dimensional optimization problem, provided that $f^{\lambda c}$ is known. The main observation of this paper, which we show in section 2, is that there exists a tractable representation of $f^{\lambda c}$ if $f$ is convex. This case is important in mathematical finance since non-negative combinations of call and put options are convex.
@@
In section 3 we use the dual problem for robust expectation to derive a representation of robust Expected Shortfall in terms of an optimization over two parameters, and apply this to derive an analytical expression for the robust Expected Shortfall of a call option, from the point of view of the writer, under the risk-neutral distribution.
@@ 2 $\lambda c$-transform for convex, piecewise-linear, functions
where $\textrm{Cpl}(\mu,\nu)$ denotes the set of “couplings” of $\mu$ and $\nu$; that is, the set of probability measures $\pi$ on $X\times X$ with first marginal distribution equal to $\mu$, and second to $\nu$. A coupling will always exist, take for example $\pi$ as the product measure $\mu\otimes\nu$. With the quadratic cost function, it holds [9, Theorem 4.1] that $d_{c}(\mu,\nu)<\infty$ if $\mu$ and $\nu$ have finite second moments; that is, if $\int_{X}\|x_{0}-x\|^{2}d\mu(x)<\infty$ and $\int_{X}\|x_{0}-x\|^{2}d\nu(x)<\infty$ for any $x_{0}\in X$.
In the case of the quadratic cost function it is well known the infimum in Equation (3) is obtained by a transport map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ that defines $\nu=T_{\#}(\mu)$, in other words the image measure induced by $\nu_{0}$ acting through $T$.
@@
As mentioned in the introduction, the robust expected value problem is tractable once the $\lambda c$-transform $f^{\lambda c}$ is available. In the case of convex $f$, the following representation of $f^{\lambda c}$ is known.
@@
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a convex function, represented as $f(x)=\sup_{i\in\mathbb{N}}\{\langle m_{i},x\rangle+c_{i}\},$ with $m_{i}\in\mathbb{R}^{d},c_{i}\in\mathbb{R},i\in\mathbb{N}.$ Let $c$ denote the cost function $c(x,y)=\frac{1}{2}\|x-y\|^{2}$. Then
$f^{\lambda c}(x)=\sup_{i\in\mathbb{N}}\left\{\langle m_{i},x\rangle+c_{i}+\frac{\|m_{i}\|^{2}}{2\lambda}\right\}.$
In particular, if $f$ is piecewise-linear and convex, then so is $f^{\lambda c}$.
Proof.
It is well known that every convex function on $\mathbb{R}^{d}$ is a pointwise supremum of a countable collection of affine functions. Let $f_{i}(x)=\langle m_{i},x\rangle+c_{i}$. Then
$f^{\lambda c}(x)$ $=$ $\sup_{y\in\mathbb{R}^{d}}\{f(y)-\lambda c(x,y)\}$
$=$ $\sup_{y\in\mathbb{R}^{d}}\{\sup_{i\in\mathbb{N}}f(y)-\lambda c(x,y)\}$
$=$ $\sup_{i\in\mathbb{N}}f^{\lambda c}_{i}(x).$
To find $f^{\lambda c}_{i}$ observe that the mapping $y\mapsto f_{i}(y)-\frac{1}{2\lambda}\|x-y\|^{2}$ is concave and differentiable, so it is sufficient to apply the first-order condition $m_{i}-\lambda(x-y)=0$; that is, $y=x+\frac{1}{\lambda}m_{i}$. So $f^{\lambda c}_{i}(x)=f_{i}(x+\frac{1}{\lambda}m_{i})=\langle m_{i},x\rangle+c_{i}+\frac{1}{2\lambda}\|m_{i}\|^{2}$. ∎
This result shows that under quadratic transportation cost, the $\lambda c$-transform shifts each affine component upwards by a constant proportional to the square of its slope. A partial heuristic interpretation is given in Figure 2, where the $\lambda c$-transforms, of payoff functions that differ by a constant $k$, also differ by $k$. This is in line with the intuition that a payoff, independent of the underlying, is not exposed to the distribution risk of the underlying.
@@
Put option with strike $K_{1}$ and two call options with strike $K_{2}>K1$. Then $f(x)=\max\{K_{1}-x,0\}+2\max\{x-K,0\}=\max\{-x+K_{1},2x-2K_{2},0\}$. Using Theorem 1,
$f^{\lambda c}(x)=\max\left\{-x+K_{1}+\frac{1}{2\lambda},2x-2K_{2}+\frac{2}{\lambda},0\right\}=\max\left\{-x+K_{1}+\frac{1}{2\lambda},2(x-(K_{2}-\frac{1}{\lambda})),0\right\}.$
@@
Figure 3: $\lambda c$-transform adjusts $f$ based on the slope of the affine component.
@@
$f^{\lambda c}(x)=\max_{i=1,\dots,4}\left\{\langle m_{i},x\rangle+c_{i}+\frac{\|m_{i}\|^{2}}{2\lambda}\right\}.$ (5)
@@ 3 Robust Expected Shortfall
We consider expected shortfall as a risk measure $\rho$ on payoff functions on $\mathbb{R}^{d}$. For risk measures on functions, see [4]. In the distributional robustification process, we associate a risk measure $\rho_{\nu}$ to each $\nu$ in the uncertainty set. In particular, Expected Shortfall of the same payoff under different probability distributions, are seen as different risk measures in the sense of giving different risk assessments.
Definition 5.
Let $\mathcal{N}$ be a set of Borel probability measures on $\mathbb{R}^{d}$ such that $\nu_{0}\in A$, and $\rho_{\nu}$ a risk measure based on $\nu$, for every $\nu\in\mathcal{N}$. Then $\tilde{\rho}(f):=\sup\{\rho_{\nu}(f):\ \nu\in\mathcal{N}\}$ is the robustified risk measure.
It may be observed that if each $\rho_{\nu}$ is a coherent risk measure, then so is $\rho$.
For simplicity we consider a liability $f$ on an atomless space, so that $\textrm{ES}_{\beta}^{\ \nu}(f)=\frac{1}{1-\beta}\int_{w}^{\infty}f(x)d\nu(x)$, where $w:=\textrm{VaR}_{\beta}^{\nu}(f)$ is the $\beta$-quantile of $f$ under the measure $\nu$, and $\beta$ the confidence level under which Expected Shortfall is computed. Let us denote the baseline risk-neutral measure by $\nu_{0}$. As in [1] we use the “minimizing property” that characterizes $\textrm{ES}_{\beta}$, see Rockafellar and Uryasev [8, Theorem 1],
@@
This expression will be referred to as the dual formula for Expected Shortfall, as it can also be derived using the dual representation of that risk measure. We recall from [8] that $\alpha^{*}:=\textrm{VaR}_{\beta}(f)$ is a minimizer in this expression. In this section $d=1$. Our uncertainty set is
@@
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be Lipschitz continuous. The robust $\textrm{ES}_{\beta}(f)$ namely
@@
$\tilde{\rho}(f)=\inf\left\{\alpha+\frac{\lambda\theta}{1-\beta}+\frac{1}{1-\beta}\int_{\mathbb{R}^{d}}((f(x)-\alpha)^{+})^{\lambda c}\ d\nu_{0}(x):\ \alpha\in\mathbb{R},\lambda\geq 0\right\}.$ (7)
@@
$=$ $\sup_{\nu\in\mathcal{N}}\min_{\alpha\in\mathbb{R}}\left\{\alpha+\frac{1}{1-\beta}\int_{\mathbb{R}^{d}}(f(x)-\alpha)^{+}\,d\nu(x)\right\}$
$=$ $\sup_{\nu\in\mathcal{N}}\min_{v_{min}\leq\alpha\leq v_{max}}\left\{\alpha+\frac{1}{1-\beta}\int_{\mathbb{R}^{d}}(f(x)-\alpha)^{+}\,d\nu(x)\right\},$
@@
To see that an upper bound $v_{max}$ exists, let $\|f\|_{Lip}$ be the Lipschitz constant of $f$ and $W_{1},W_{2}=\sqrt{d_{c}}$ denote the respective p-Wasserstein distances. It follows from the Markov inequality, 1-Wasserstein duality [9, Remark 6.5], and the basic property $W_{1}\leq W_{2}$ [9, Remark 6.6] that for $t>0$ ,
$\nu(\{x:f(x)>t\})$ $\leq$ $\frac{1}{t}\int_{\mathbb{R}^{d}}|f(x)|d\nu(x)$
$\leq$ $\frac{1}{t}\left(\int_{\mathbb{R}^{d}}|f(x)|d\nu_{0}(x)+\|f\|_{Lip}W_{1}(\nu,\nu_{0})\right)$
$\leq$ $\frac{1}{t}\left(\int_{\mathbb{R}^{d}}|f(x)|d\nu_{0}(x)+\|f\|_{Lip}\sqrt{\theta}\right)$
@@
Since $\alpha$ ranges over a compact set, and the function $(\alpha,\nu)\mapsto\alpha+\frac{1}{1-\beta}\int_{\mathbb{R}^{d}}(f(x)-\alpha)^{+}\,d\nu(x)$, being convex in $\alpha$ and linear in $\nu$, is easily seen to be “convex-concavelike” [3], we may use Ky Fan’s minimax theorem. Thus
$\tilde{\rho}(f)$ $=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\sup_{\nu\in\mathcal{N}}\left\{\alpha+\frac{1}{1-\beta}\int_{\mathbb{R}^{d}}(f(x)-\alpha)^{+}\,d\nu(x)\right\}$
$=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\sup_{\nu\in\mathcal{N}}\int_{\mathbb{R}^{d}}\alpha+\frac{1}{1-\beta}(f(x)-\alpha)^{+}d\nu(x)$
$=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\left\{\alpha+\frac{1}{1-\beta}\sup_{\nu\in\mathcal{N}}\int_{\mathbb{R}^{d}}(f(x)-\alpha)^{+}d\nu(x)\right\}.$
Applying robust duality Equation (2) to the function $x\mapsto(f(x)-\alpha)^{+}$, we get
$\tilde{\rho}(f)$ $=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\left\{\alpha+\frac{1}{1-\beta}\inf_{\lambda\geq 0}\left[\lambda\theta+\int_{\mathbb{R}^{d}}((f(x)-\alpha)^{+})^{\lambda c}d\nu_{0}(x)\right]\right\}$
$=$ $\min_{v_{min}\leq\alpha\leq v_{max}}\inf_{\lambda\geq 0}\left\{\alpha+\frac{1}{1-\beta}\lambda\theta+\frac{1}{1-\beta}\int_{\mathbb{R}^{d}}((f(x)-\alpha)^{+})^{\lambda c}d\nu_{0}(x)\right\}$
Writing the minimum as an infimum, and combining the infimums, we get the result. ∎
We demonstrate Theorem 6 with a simple toy example.
Example 7.
Consider a uniform distribution $\nu_{0}$ on the unit interval $[0,1]$, and let $f(x)=x$ for $x\in\mathbb{R}$. The minimum in Equation (2) can be obtained in closed form. The first-order conditions for $\alpha$ and $\lambda$ yield the solutions $\alpha=\beta+\frac{1}{2\lambda}$ and $\lambda=\sqrt{\frac{1-\beta}{2\theta}}$. This corresponds to the transport map
$T_{\lambda}(x)=\begin{cases}x+\frac{1}{\lambda}&\text{ if $\beta<x<1$}\\ x&\text{ if $0<x\leq\beta$}\end{cases},$
and we see that the constraint $d_{c}(\nu_{0},\nu)\leq\theta$ is satisfied, where $\nu=(T_{\lambda})_{\#}(\nu_{0})$, since $\int_{0}^{1}\frac{1}{2}|T_{\lambda}(x)-x|^{2}dx=\frac{1}{2}\int_{\beta}^{1}(\frac{1}{\lambda})^{2}dx=\theta$. As could be expected, the robustification procedure shifts the upper tail of the distribution as far as possible while respecting the transport cost constraint.
@@
Since $\nu_{0}$ is a pricing measure, $\int g^{\lambda c}(x)\,d\nu_{0}(x)=\textrm{call}(k+\alpha-\frac{1}{2\lambda})$. By Theorem 6,
$\tilde{\rho}(f)=\min_{v_{min}\leq\alpha\leq v_{max}}\inf_{\lambda\geq 0}\left\{\alpha+\frac{1}{1-\beta}\left[\lambda\theta+\textrm{call}(k+\alpha-\frac{1}{2\lambda})\right]\right\}.$ (8)
@@
In contrast to robustifications of ES based on the distribution of $f(X)$, [10][1, Table 1], the “robustification correction” $\tilde{\rho}(f)-\rho_{\nu_{0}}(f)$ term for Expected Shortfall is a function of $f$, since it depends on $k$. Indeed,
$\tilde{\rho}(f)=\rho_{\nu_{0}}(f)+\sqrt{\frac{2\theta}{1-\beta}}+\begin{cases}q_{\beta}-k&\text{ if }k\leq q_{\beta},\\ 0&\text{ if }k>q_{\beta}.\end{cases}$ (10)
This difference is due to the definition of the risk measure in terms of the payoff $f$ instead of the distribution, and the fact that in the ES maximization process, imagining an adversary that wants to maximize risk, it is only worthwhile to transport mass that is already on the right hand side of the strike price.
In conclusion, we see that robustification adds a term proportional to the square root of the ambiguity radius, and inversely proportional to $1-\beta$. In particular, this quantifies the high distributional sensitivity of an Expected Shortfall calculation for a confidence level $\beta\approx 1$.